OpenStudy (anonymous):

What is the distance between points (6, 5) and (–12, 17)? Round to the nearest tenth. Answer A. 8.9 B. 14.7 C. 18.2 D. 21.6

OpenStudy (jamesj):

Distance is the square root, sqrt, of change in x squared plus change in y squared. I.e., distance^2 = (6-(-12))^2 + (5-17)^2 = 18^2 + (-12)^2 = 324 + 144 = 468 So distance = ...

OpenStudy (anonymous):

soooo its C?